### v2

# Cooperative and Asynchronous Transformer-based Mission Planning for Heterogeneous Teams of Mobile Robots

Abstract
Cooperative mission planning for heterogeneous teams of mobile robots presents a unique set of challenges, particularly when operating under communication constraints and limited computational resources. To address these challenges, we propose the Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework, which leverages multi-agent reinforcement learning (MARL) to coordinate distributed decision making among agents with diverse sensing, motion, and actuation capabilities, operating under sporadic ad hoc communication. A Class-based Macro-Action Decentralized Partially Observable Markov Decision Process (CMacDec-POMDP) is also formulated to effectively model asynchronous decision-making for heterogeneous teams of agents. The framework utilizes an asynchronous centralized training and distributed execution scheme that is developed based on the Multi-Agent Transformer (MAT) architecture. This design allows a single trained model to generalize to larger environments and accommodate varying team sizes and compositions. We evaluate CATMiP in a 2D grid-world simulation environment and compare its performance against planning-based exploration methods. Results demonstrate CATMiP’s superior efficiency, scalability, and robustness to communication dropouts, highlighting its potential for real-world heterogeneous mobile robot systems. The code is available at https://github.com/mylad13/CATMiP.
keywords:
Autonomous Robots , Motion Planning , Robot Learning , Reinforcement Learning.
journal: Robotics and Autonomous Systems
[…]
organization=Department of Electrical and Computer Engineering, McMaster University,addressline=1280 Main St W, city=Hamilton, postcode=L8S 4L8, state=ON, country=Canada
1 INTRODUCTION
Multi-robot systems (MRS) are becoming increasingly prevalent in applications such as search and rescue operations [1], environmental monitoring [2], building and infrastructure inspection, [3], and industrial plant management [4]. The coordinated efforts of robots in these systems improve efficiency and adaptability, especially in complex tasks. Particularly, heterogeneous MRS composed of robots with complimentary capabilities outperform homogeneous teams in missions requiring diverse sensing and actuation capabilities [5, 6, 7].
Coordination in MRS can be centralized or decentralized. Centralized approaches rely on a leader robot or server to issue commands, which can result in high computational and communication loads, vulnerability to single-point failures, and challenges in ensuring consistent communication in real-world scenarios. In contrast, decentralized approaches allow robots to individually make autonomous decisions while implicitly considering the actions of others and changes in the environment. Combining implicit coordination with explicit communication through ad hoc wireless mesh networks enables distributed control strategies that are both scalable and efficient [8, 9].
Deep Multi-Agent Reinforcement Learning (MARL) has emerged as a powerful tool for coordinating MRS in dynamic and uncertain environments [10]. Deep MARL enables robots to learn coordination strategies autonomously, bypassing the need for predefined algorithms and heuristics. However, conventional MARL often assumes synchronous decision-making, where agents take new actions at the same time—a condition that is inefficient and impractical for many real-world scenarios. Asynchronous MARL [11] addresses this limitation by enabling agents to make decisions over temporally extended actions, otherwise known as macro-actions [12].
This paper addresses the distributed coordination of heterogeneous mobile robots navigating unknown environments by proposing the Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework. CATMiP is formulated based on the Class-based Macro-Action Decentralized Partially Observable Markov Decision Process (CMacDec-POMDP) model, a novel extension of the MacDec-POMDP model [12] for decentralized multi-agent planning that considers varying properties across different agent classes. Our case study involves two robot types-explorers and rescuers, where the objective is for a rescuer type robot to reach a target with an initially unknown location as fast as possible.
The robots perform collaborative simultaneous localization and mapping (C-SLAM), merging local occupancy grid maps into a shared global map through intermittent communication [13]. Navigation decisions are made in a distributed manner and following a hierarchical two-level control approach. At the high-level, each robot selects a macro-action, which is a goal point location on the map within a fixed distance from the robot. This macro-action is sampled from a policy generated by the proposed Asynchronous Multi-Agent Transformer (AMAT) network. The inputs to AMAT are the agents’ macro-observations, which are class-specific multi-channeled global and local maps. At the low-level, path planning and motion control modules generate the robot’s immediate action to navigate toward the selected goal.
Section 2 reviews related works and discusses how the identified research gaps are addressed in this paper. Section 3 introduces the CMacDec-POMDP formulation, and formally states the problem. Section 4 details the different components of the CATMiP framework, as well as the asynchronous training and execution phases of AMAT. Section 6 describes the simulation setup, including macro-observations, macro-actions, and reward structures. Simulation results are presented and analyzed in Section 7. Finally, Section 8 concludes the paper and outlines future directions.
2 Related Works
Deep reinforcement learning (DRL) has been increasingly used in mobile robotics for exploration and navigation, handling complex tasks in single-agent and multi-agent settings [14, 10, 15]. Such control strategies are typically divided into end-to-end and two-stage approaches. The end-to-end methods derive control actions directly from sensor data, whereas the two-stage approaches first select target locations using DRL and then employ a separate method for control actions, improving sample efficiency and generalization. Notable recent works have combined high-level DRL-based goal selection with classical path-planning algorithms in single robot scenarios [16, 17, 18].
Cooperative multi-robot mission planning has been studied using various deep MARL approaches. Notable works addressing asynchronous multi-robot exploration with homogeneous robots and macro-actions include [19, 20]. Tan et al. [19] tackle the challenge of communication dropouts in multi-robot exploration by modeling the problem as a MacDec-POMDP and proposing a DRL solution based on the centralized training and decentralized execution (CTDE) paradigm [21]. CTDE strikes a balance between coordination and scalability by enabling agents to learn from shared experiences during training while acting independently based on local observations during execution. Yu et al. [20] extend the multi-agent proximal policy optimization (MAPPO) algorithm [22] to enable asynchronous CTDE. Their approach enhances coordination efficiency through an attention-based relation encoder, which aggregates feature maps from different agents to capture intra-agent interactions. While these methods demonstrate the effectiveness of macro-actions in asynchronous decision-making, all agents follow the same policy and the unique challenges of planning for heterogeneous multi-robot systems are not considered.
To address heterogeneity, Zhang et al. [23] propose an architecture for asynchronous multi-robot decision-making that combines value function decomposition [24], the MacDec-POMDP framework, and the CTDE paradigm. Their approach utilizes features extracted from both global states and local observations during training. However, during execution, each agent generates macro-actions based solely on feature maps derived from its local observations. This design enables diverse behaviors among agents but restricts the trained model to be used by a fixed team size and composition. Moreover, none of the aforementioned methods allow the trained models to generalize to larger environments. This limitation highlights the need for approaches that prioritize scalability and adaptability in multi-robot mission planning.
To enable agent heterogeneity while maintaining the benefits of parameter sharing, agent indication was formalized in [25]. This method appends an agent-specific indicator signal to the observations, allowing a shared policy network to generate agent-specific actions. Terry et al. [25] demonstrated that parameter sharing can be effectively applied to heterogeneous observation and action spaces while still achieving optimal policies. This idea is used in the Multi-Agent Transformer (MAT) architecture [26] as well, where positional encoding that appears in the original transformer [27] are replaced by agent indication.
Wen et al. [26] introduced MAT alongside a novel MARL training paradigm that achieves linear time complexity and guarantees monotonic performance improvement by leveraging the multi-agent advantage decomposition theorem [28]. This theorem suggests that joint positive advantage can be achieved by sequentially selecting local actions rather than searching the entire joint action space simultaneously. Thus, cooperative MARL can be reformulated as a sequence modeling problem, where the objective is to map a sequence of agent observations to a sequence of optimal agent actions.
In MAT, the attention mechanism [27] in the encoder captures the inter-agent relationships within the sequence of observations, and the decoder auto-regressively generates actions by considering the input sequence’s latent representation. The transformer model’s ability to process flexible sequence lengths enables generalization to different team sizes without treating varying agent numbers as separate tasks. This property allows a single trained model to scale to teams with more or less agents than those encountered during training.
Building on MAT, our work introduces the Asynchronous Multi-Agent Transformer (AMAT) network. We develop a new asynchronous centralized training and asynchronous distributed execution scheme tailored for heterogeneous teams. Specialized agent class policies are learned through agent class encodings, which differentiate macro-observations across agent classes and enable the network to generate corresponding macro-actions. This design enhances the model’s generalizability to larger teams with varying compositions of heterogeneous agents. Additionally, we employ an adaptive pooling layer during global feature extraction from macro-observations, allowing the model to efficiently scale to larger environment sizes without compromising performance.
3 Problem Statement and Formulation
This paper addresses the design of distributed controllers for a heterogeneous team of mobile robots performing a cooperative mission in an unknown environment. Specifically, we focus on an indoor search and target acquisition scenario involving two agent classes: explorers and rescuers. The mission objective is for a rescuer agent to reach a target with an initially unknown location as quickly as possible. To achieve this, control policies must exploit the diverse capabilities of the team, encouraging specialized behaviors for each agent class. For instance, explorer robots, being faster and more agile, are tasked with rapidly mapping the environment and locating the target. In contrast, rescuer robots, though slower, have the capability to engage with the target once its location is known.
To solve this problem, we propose a hierarchical control approach that combines high-level decision-making and low-level motion control for effective navigation. First, a goal location within a localized area centered on the robot is selected by the high-level decision-making module; then, the local planner generates motion commands to ensure smooth movement and obstacle avoidance en route to the goal.
We formalize this approach as a Class-based Macro Action Decentralized Partially Observable Markov Decision Process (CMacDec-POMDP), a novel extension of the MacDec-POMDP framework [12] that incorporates varying properties across different agent classes. Assuming a team of $N$ agents with $C\leq N$ different classes, the problem is formalized as the tuple $\langle I,C,S,\{M^{c}\},\{A^{c}\},T,\{R^{c}\},\allowbreak\{\zeta^{c}\},\{Z^{c}\},\{\Omega^{c}\},\{O^{c}\},\gamma,h\rangle$, where
  • 1.
    $I=\{1,\dots,N\}$ is a finite set of agents;
  • 2.
    $C=\{1,\dots,C\}$ is a finite set of agent classes, with $C(i)$ indicating the class of agent $i$;
  • 3.
    $S$ is the global state space;
  • 4.
    $M^{c}$ is a finite set of macro-actions (MAs) for agents of class $c$. The set of joint MAs is then $M=\times_{i}M^{C(i)}$;
  • 5.
    $A^{c}$ is a finite set of (primitive) actions for agents of class $c$. The set of joint actions is then $A=\times_{i}A^{C(i)}$;
  • 6.
    $T:S\times A\times S\rightarrow[0,1]$ is a state transition probability function, indicating the probability of transitioning from state $s\in S$ to state $s^{\prime}\in S$ when the agents are taking the joint action $\bar{a}\in A$. In other words, $T(s,\bar{a},s^{\prime})=Pr(s^{\prime}|\bar{a},s)$;
  • 7.
    $R^{c}:S\times A\rightarrow\mathbb{R}$ is the agent-class-specific reward function, with $R^{C(i)}(s,\bar{a})$ being the reward an agent $i$ of class $c$ receives when the joint action $\bar{a}$ is executed in state $s$;
  • 8.
    $\zeta^{c}$ is a finite set of macro-observations (MOs) for agents of class $c$. The set of MOs is then $\zeta=\times_{i}\zeta^{C(i)}$;
  • 9.
    $Z^{c}:\zeta^{c}\times M^{c}\times S\rightarrow[0,1]$ is the MO probability function for agents of class $c$, indicating the probability of the agent receiving the MO $z^{i}\in\zeta^{C(i)}$ given MA $m^{i}\in M^{C(i)}$ is in progress or has completed and the current state is $s^{\prime}\in S$. In other words, $Z^{C(i)}(z^{i},m^{i},s^{\prime})=Pr(z^{i}|m^{i},s^{\prime})$;
  • 10.
    $\Omega^{c}$ is a finite set of observations for agents of class $c$. The set of joint observations is then $\Omega=\times_{i}\Omega^{C(i)}$;
  • 11.
    $O^{c}:\Omega^{c}\times A^{c}\times S\rightarrow[0,1]$ is the observation probability function for agents of class $c$, indicating the probability of agent $i$ receiving the observation $o^{i}\in\Omega^{C(i)}$ when the current state is $s^{\prime}\in S$ after the agents have taken the joint action $\bar{a}\in A$. In other words, $O^{C(i)}(o^{i},\bar{a},s^{\prime})=Pr(o^{i}|\bar{a},s^{\prime})$;
  • 12.
    $\gamma\in[0,1]$ is the discount factor;
  • 13.
    and $h$ is the horizon, the number of steps in each episode.
Throughout the mission, each agent $i$ selects an MA $m^{i}$ using the high-level policy $\mu^{i}:H_{M}^{i}\times M^{C(i)}\rightarrow[0,1]$. An MA is represented as $m^{i}=\langle\beta_{m^{i}},\mathcal{I}_{m^{i}},\pi_{m^{i}}\rangle$, where $\beta_{m^{i}}:H_{A}^{i}\rightarrow[0,1]$ is a stochastic termination condition based on the primitive action-observation history $H_{A}^{i}$, $\mathcal{I}_{m^{i}}\subset H^{i}_{M}$ is the initiation condition that determines whether the MA can be started based on the macro-action-macro-observation history $H^{i}_{M}$, and $\pi_{m^{i}}:H^{i}_{A}\times A^{C(i)}\rightarrow[0,1]$ is the low-level policy that generates the primitive actions $a^{i}$ required to execute the MA.
The underlying Dec-POMDP is used to generate primitive transitions and rewards, but the low-level policy $\pi_{m^{i}}$ of agent $i$ is determined by the MA obtained via the high-level policy $\mu^{i}$. Access to the full model of the underlying Dec-POMDP is not necessary, as the MAs are assumed to be simulated in an environment close enough to the real-world domain. This allows all evaluations to be conducted in the simulator through sampling [12].
The solution to the CMacDec-POMDP is the joint high-level policy $\bar{\mu}=(\mu^{1},\dots,\mu^{N})$ that maximizes the expected cumulative discounted return from all agents,
$\bar{\mu}^{*}=\operatorname*{arg\,max}_{\bar{\mu}}~{}~{}\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=0}^{h-1}\gamma^{t}R^{C(i)}(s_{t},\bar{a}_{t})|s_{0},\bar{\pi},\bar{\mu}\right].$ (1)
It should be noted that a generic parameter $p$ is denoted as $p^{i}$ when referring to the $i$-th agent, as $p^{i_{j:k}}=(p^{i_{j}},p^{i_{j+1}},\dots,p^{i_{k}})$ when referring to a subset $i_{j:k}$ of agents, and as $\bar{p}=(p^{1},p^{2},\dots,p^{N})$ when referring to the team of $N$ agents.
Figure 1: Workflow of the CATMiP framework during the execution phase. The robots communicate to share mission information and their embedded macro-observations to make high-level navigation decisions in a distributed manner, while a local motion controller generates the immediate action for each robot.
4 System Architecture
The overall architecture of the proposed Cooperative and Asynchronous Transformer-based Mission Planning (CATMiP) framework is shown in Figure 1. CATMiP is composed of three key modules: the C-SLAM module, which provides global mapping and localization; the High-Level Decision Making module, which determines macro-actions for agents and handles inter-agent communication; and the Low-Level Motion Control module, which translates macro-actions into motion commands for navigation. To enable efficient and scalable operation, CATMiP employs an asynchronous centralized training process to learn agent policies and an asynchronous distributed execution scheme to deploy these policies during real-time missions. The architecture is designed to leverage the diverse capabilities of heterogeneous agents while maintaining scalability to larger team sizes and environments.
4.1 Collaborative Simultaneous Localization and Mapping (C-SLAM)
This module enables robots to collaboratively build a shared global occupancy grid map. A distributed C-SLAM method that works under intermittent communication conditions, such as Swarm-SLAM [13], can be utilized during the execution phase. A designated robot temporarily processes data received from connected robots to generate the most accurate map and pose estimates, ensuring convergence to a unified global reference frame after a few rendezvous among subsets of robots. The occupancy grid map is then passed unto the high-level decision-making and low-level motion control modules.
4.2 High-Level Decision-Making
The High-Level Decision Making module determines macro-actions for each agent and enables coordination by leveraging information from robot sensory data, the C-SLAM module, and inter-agent communication.
During the centralized training phase, the Mission Coordinator sub-module aggregates mission-related information and forms all agents’ MOs, denoted as $\bar{z}$. The Mission Coordinator maintains the action-observation history $\bar{H}_{A}$ and the macro-action-macro-observation history $\bar{H}_{M}$. It manages agent activation and determines new MAs by passing $\bar{z}$ through the Asynchronous Multi-Agent Transformer (AMAT) network. The selected MAs are then sent to the Low-Level Motion Control module as navigation goals for execution.
At each time step of the mission, whether during training or execution, agents are categorized as active, inactive, or on-standby, based on their current state and macro-action progress. Inactive agents continue executing primitive actions based on their current MA, as dictated by their low-level policy. Agents enter standby mode when their current MA termination condition is met. During training, agents on standby delay selecting their next MA for a random minimum number of time steps. This delay serves two purposes: it simulates the variable time required to complete MAs in real-world deployments and allows multiple standby agents to simultaneously activate and update their MAs with a single pass through the AMAT network. If an agent’s maximum standby time is reached and no other agents are on-standby, it activates alone. Agents that receive new MAs return to the inactive state to begin executing the newly assigned goal.
In the distributed execution phase, the Mission Coordinator operates locally on each robot, managing agent activation independently. It also facilitates the exchange of MOs and MAs between communicating agents. This localized approach ensures efficient real-time decision-making while maintaining coordination across the team.
4.3 Low-Level Motion Control
The Low-Level Motion Control module generates the robot’s primitive action $a^{i}_{t}$ at each time step $t$ by integrating a path planner and a motion controller. The path planner determines a collision-free path to the global goal specified by the robot’s current MA $m^{i}$. A path-finding algorithm, such as A* search [29], can be used to compute the shortest path on the occupancy grid map, represented as a sequence of waypoints.
To refine the planned path, the motion controller employs a local planner [30], which optimizes the robot’s trajectory based on the selected path and real-time sensory data. The local planner ensures that the resulting trajectory adheres to the robot’s motion constraints, avoids dynamic and static obstacles, and minimizes execution time. The optimized trajectory is then translated into low-level motion control commands, producing the primitive action $a^{i}_{t}$.
4.4 Asynchronous Centralized Training
At every time step $t$ of a training episode, the agents execute the joint primitive action $\bar{a}_{t}$ generated by their low-level motion control module, and each collect a class specific state-action-dependent reward, jointly denoted as $\bar{R}(s_{t},\bar{a}_{t})$. At the $\tau$th instance of there being a subset $i_{1:g(\tau)}$ of agents ready to update their MAs (flagged as active agents), an ordered sequence of MOs $\bar{z}_{\tau}=\{z^{i_{1}}_{\tau},\dots,z^{i_{g(\tau)}}_{\tau},z^{i_{g(\tau)+1}}_{\tau},\dots,z^{i_{N}}_{\tau}\}$ are formed and fed as the input to the AMAT network. Subsequently, a sequence of latent representation of MOs of active agents, denoted as $\hat{z}^{i_{1:g(\tau)}}_{\tau}$, are calculated and used to obtain new MAs $m^{i_{1:g(\tau)}}_{\tau}$ and value function estimates $V(\hat{z}^{i_{1:g(\tau)}}_{\tau})$. The rewards accumulated by the active agents since their last activation, $R^{i_{1:g(\tau)}}_{acc}$, along with $\bar{z}_{\tau}$, $m^{i_{1:g(\tau)}}_{\tau}$, and $V(\hat{z}^{i_{1:g(\tau)}}_{\tau})$ are stored in a replay buffer. At the end of each training episode, the stored transitions are used to optimize the AMAT network parameters through backpropagation. The pseudocode of the centralized training process is presented in Algorithm 1.
The resulting fully trained network is then deployed for decentralized or distributed execution on each robot’s local hardware, enabling real-time inference of macro-actions during the mission.
Algorithm 1 Centralized Training of AMAT
1:Number of agents $n$, episodes $K$, steps per episode $h$, minibatch size $B$.
2:Encoder $\{\phi_{0}\}$, Decoder $\{\theta_{0}\}$, Replay Buffer $\mathcal{B}$.
3:for $k=0,1,\dots,K-1$ do
4:    Initialize $\tau=1$, $\bar{R}_{acc}=0$, and obtain $\bar{m}=\bar{m}_{0}$ and $V_{\phi_{0}}(\bar{z}_{0})$ from initial joint macro-observations $\bar{z}_{0}$ and insert $(\bar{m}_{0},\bar{z}_{0}),V_{\phi_{0}}(\bar{z}_{0})$ into $\mathcal{B}$.
5:    for $t=0,1,\dots,h-1$ do
6:        Execute joint actions $\bar{a}_{t}$ sampled from joint policies $\pi_{\bar{m}}$, and collect the joint reward $\bar{R}(s_{t},\bar{a}_{t})$.
7:        $\bar{R}_{acc}\leftarrow\bar{R}_{acc}+\bar{R}(s_{t},\bar{a}_{t})$.
8:        if non-empty subset $i_{1:g(\tau)}$ of agents are active then
9:           Form MOs $\bar{z}_{\tau}$, obtain MO embeddings $\bar{z}_{e,\tau}$ and feed them to the encoder to generate the latent representations of MOs for the active subset of agents, $\hat{z}_{\tau}^{i_{1:g(\tau)}}$.
10:           Generate $V_{\phi_{k}}(\hat{z}_{\tau}^{i_{1}}),\dots,V_{\phi_{k}}(\hat{z}_{\tau}^{i_{g(\tau)}})$ with the output layer of the encoder.
11:           Input $\hat{z}_{\tau}^{i_{1:g(\tau)}}$ to the decoder.
12:           for $l=0,1,\dots,g(\tau)-1$ do
13:               Input $m_{\tau}^{i_{0}},\dots,m_{\tau}^{i_{l}}$ and infer $m_{\tau}^{i_{l+1}}$ with the auto-regressive decoder.
14:           end for
15:           Update the MAs of active agents $\{m^{i_{1}},\dots,m^{i_{g(\tau)}}\}\leftarrow\{m_{\tau}^{i_{1}},\dots,m_{\tau}^{i_{g(\tau)}}\}$.
16:           Set $R_{\tau-1}^{i_{1:g(\tau)}}\leftarrow R_{acc}^{i_{1:g(\tau)}}$.
17:           Insert $(\bar{z}_{\tau},m_{\tau}^{i_{1:g(\tau)}},R_{\tau-1}^{i_{1:g(\tau)}},V_{\phi_{k}}(\hat{z}_{\tau}^{i_{1:g(\tau)}}))$ into $\mathcal{B}$.
18:           Set $R_{acc}^{i_{1:g(\tau)}}\leftarrow 0$.
19:           Set $\tau\leftarrow\tau+1$.
20:        end if
21:    end for
22:    Set $T\leftarrow\tau$ as the number of experiences.
23:    Sample a random minibatch of $B$ experiences from $\mathcal{B}$.
24:    Compute the advantage function with GAE using the value function estimates.
25:    Minimize $L_{Encoder}(\phi)+L_{Decoder}(\theta)$ (Equations 3,4) on the minibatch with gradient descent and update the encoder and decoder to obtain $\phi_{k+1}$ and $\theta_{k+1}$
26:end for
4.5 Asynchronous Distributed Execution
During the execution phase, the robots operate in a fully distributed manner, utilizing local instances of the AMAT network and time-varying communication neighborhoods formed through a dynamic mobile ad hoc network [31]. The communication topology is modeled as a dynamic graph $\mathcal{G}=(\mathcal{I},\mathcal{E})$, where vertices $\mathcal{I}$ represent the set of agents, and edges $\mathcal{E}$ represent communication links between them. The probability of a communication link $E_{ij}$ existing between agents $i$ and $j$ depends on their distance $d_{ij}$, defined as:
$Pr(E_{ij})=e^{-d_{ij}^{2}/\sigma^{2}},$ (2)
where $\sigma$ is a decay parameter controlling how quickly the communication probability decreases with distance. This probabilistic model is adapted from [32]. Each agent $i$ maintains a local communication neighborhood $\mathcal{N}(i)$, consisting of all agents (including itself) that can exchange information either directly or indirectly at a given time step.
Agents within the same communication neighborhood that become active at the same time step coordinate to elect a temporary broker. The broker facilitates local decision-making by aggregating neighborhood information and generating macro-actions for all active agents in its vicinity. It forms an ordered sequence of MO embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{g}},z_{e}^{i_{g+1}},\dots,z_{e}^{i_{m}})$, where $z_{e}^{i_{1}}$ corresponds to the broker’s own MO embedding, $z_{e}^{i_{2:g}}$ are embeddings received from other active agents in the neighborhood, and $z_{e}^{i_{{g+1}:m}}$ are the current MO embeddings from inactive agents in the neighborhood. The broker feeds this ordered sequence into its local instance of the AMAT network. The network outputs a sequence of updated MAs $(m^{i_{1}},\dots,m^{i_{g}})$ for all active agents in the neighborhood. The broker transmits the newly generated macro-actions back to the respective agents. These macro-actions serve as navigation goals and are passed to each robot’s local Low-Level Motion Control module to execute corresponding primitive actions.
5 Asynchronous Multi-Agent Transformer (AMAT)
This section details the structure and different components of the AMAT network and its use during asynchronous centralized training and distributed execution.
The AMAT network, illustrated in Fig. 2, consists of four components: Macro-Observations Embedder, Encoder, Macro-Actions Embedder, and Decoder. AMAT transforms a sequence of MOs $z^{i_{1:m}}$ from a subset $i_{1:m}$ of agents into a sequence of MAs $m^{i_{1:g}}$ corresponding to the active subset $i_{1:g}$ ($g\leq m$).
During centralized training, $i_{1:m}$ represents the complete agent set $I=\{1,\dots,N\}$, while in distributed execution, it refers to the agents in the neighborhood $\mathcal{N}(i)=\{i_{1},\dots,i_{m}\}$. In both cases, the sets are ordered to put the active agents $i_{1:g}$ first.
Figure 2: Centralized macro-action inference as a part of the training process of AMAT. During distributed execution, the broker robot receives macro-observation embeddings from connected agents and transmits the newly obtained macro-actions back to them.
5.1 Macro-Observations Embedder
During centralized training, this module transforms a sequence of MOs $(z^{i_{1}},\dots,z^{i_{m}})$ into embeddings $(z_{e}^{i_{1}},\dots,z_{e}^{i_{m}})$ as input tokens for the encoder. In the distributed execution phase, each agent $i_{l}\in\mathcal{N}(i)$ locally processes its MO $z^{i_{l}}$ using the trained embedder, then transmits the resulting embedding $z_{e}^{i_{l}}\in\mathbb{R}^{d}$ to its neighborhood’s broker. This approach reduces network traffic by transmitting compact embeddings instead of high-dimensional MOs, which include local and global maps.
The architecture, shown in Fig. 3, uses two separate convolutional neural networks (CNNs) to extract features from multi-channeled global and local maps. An adaptive max pooling layer scales global feature maps of any $S\times S$ dimensions to fixed $G\times G$ zones, enabling the model’s operation in varying environment sizes. The features are then concatenated and processed by a multi-layer perceptron (MLP).
Similar to the idea of agent indication [25], a learnable agent class encoding is used to enable unique behaviors for different classes of agents. This class encoding is obtained by processing a one-hot agent class identifier tied to each agent’s MO $z^{i_{l}}$ through a fully connected layer (FC). The outputs of the MLP and FC layers are combined via element-wise addition, resulting in the macro-observation embedding $z^{i_{l}}_{e}$.
Figure 3: The macro-observation embedder network, where agent $i_{l}$’s macro-observation $z^{i_{l}}$ is transformed into a macro-observation embedding $z_{e}^{i_{l}}\in\mathbb{R}^{d}$
5.2 Encoder
The encoder is made up of several encoding blocks each consisting of a self-attention mechanism, an MLP, and residual connections. It processes the sequence of MO embeddings $z_{e}^{i_{1:m}}$ into a sequence of MO representations $(\hat{z}^{i_{1}},\dots,\hat{z}^{i_{m}})$, which carry information both about each agent’s current view of the environment, as well as the high-level interrelationships among the agents. An additional MLP is also used during the training phase to approximate the value of each agent’s observation. Values associated with the active subset of agents, $(V_{\phi}(\hat{z}^{i_{1}}),\dots,V_{\phi}(\hat{z}^{i_{g}}))$, are used to minimize the empirical Bellman error
$L_{Encoder}(\phi)=\frac{1}{T}\sum_{\tau=0}^{T-1}\frac{1}{g(\tau)}\sum_{l=1}^{g(\tau)}\Bigl{[}R_{\tau}^{i_{l}}+\gamma V_{\bar{\phi}}(\hat{z}_{\tau+1}^{i_{l}})-V_{\phi}(\hat{z}_{\tau}^{i_{l}})\Bigr{]},$ (3)
where $T$ is the total number of MA updates, $g(\tau)$ is the number of active agents at the $\tau$th MA update instance, $\phi$ represents MO-embedder and encoder parameters, and $\bar{\phi}$ represents the non-differentiable target network’s parameters.
5.3 Macro-Actions Embedder
This module converts one-hot encoded representation of MAs $m^{i_{0:l-1}},l=\{1,\dots,m\}$ into MA embeddings $m_{e}^{i_{0:l-1}}$ using an MLP. Similar to the MO embedder, class encodings are combined with these embeddings to associate each MA with the agent class responsible for executing it.
5.4 Decoder
The decoder processes the joint MA embeddings $m_{e}^{i_{0:l-1}}$, $l=\{1,\dots,m\}$ through a series of decoding blocks, with $m_{e}^{i_{0}}$ acting as an arbitrary token designating the start of decoding. Each decoding block is made up of a masked self-attention mechanism, a masked attention mechanism, and an MLP followed by residual connections. The masking ensures that each agent is only attending to itself and the agents preceding it, preserving the sequential updating scheme and the monotonic performance improvement guaranty during training [26]. The final decoder block outputs a sequence of joint MA representations $\{\hat{m}^{i_{0:h-1}}\}_{h=1}^{l}$, which is then fed to an MLP to obtain the probability distribution of agent $i_{l}$’s MA, which is the high-level policy $\mu^{i_{l}}_{\theta}(m^{i_{l}}|\hat{z}^{i_{1:m}},\hat{m}^{i_{0:l-1}})$, where $\theta$ represents the MA-embedder and decoder parameters. The decoder is trained by minimizing the following clipped PPO objective, which only uses the action probabilities and advantage estimates of the active subset of agents $i_{1:g(\tau)}$ at the $\tau$th instance of MA updates.
$L_{Decoder}(\theta)=-\frac{1}{T}\sum_{\tau=0}^{T-1}\frac{1}{g(\tau)}\sum_{l=1}^{g(\tau)}\min(r_{\tau}^{i_{l}}(\theta)\hat{A}_{\tau}^{i_{l}},clip(r_{\tau}^{i_{l}}(\theta),1\pm\epsilon)\hat{A}_{\tau}^{i_{l}}),$ (4)
$r_{\tau}^{i_{l}}(\theta)=\frac{\mu^{i_{l}}_{\theta}(m_{\tau}^{i_{l}}|\hat{z}_{\tau}^{i_{1:g}},m_{\tau}^{i_{0:l-1}})}{\mu^{i_{l}}_{\theta_{old}}(m_{\tau}^{i_{l}}|\hat{z}_{\tau}^{i_{1:g}},m_{\tau}^{i_{0:l-1}})},$ (5)
where $\hat{A}_{\tau}^{i_{l}}$ is the estimate of agent $i_{l}$’s advantage function obtained using generalized advantage estimation (GAE) [33], with $V(\hat{z}^{i_{l}}_{\tau})$ used as the estimate for the value function.
It should be noted that actions are generated in an auto-regressive manner during the inference stage, which means that generating $m^{i_{l+1}}$ requires $m^{i_{l}}$ to be inserted back into the decoder. However, the output probability of all MAs $m^{i_{1:m}}$ can be computed in parallel during the training stage since $m^{i_{1:m-1}}$ have already been collected and saved in the replay buffer. Since the attention computations are still masked, the tokens representing the MAs of the inactive subset of agents do not impact the calculations for the active subset, and can simply be replaced by zero padding.
6 Simulation Setup
6.1 2D Simulation Environment
CATMiP is evaluated in a customized 2D grid-world environment built using Minigrid [34]. Each episode features randomly generated environments of size $S\times S$, consisting of cluttered rooms with diverse shapes and a static target placed at an unknown location.
A snapshot of the simulation is shown in Fig. 4, where a rescuer robot (red triangle) and two explorer robots (blue and yellow triangles) navigate the environment to locate the target (green square). Each robot’s global goal is marked by a corresponding colored circle. Cells are either free space (black) or occupied by walls or objects (orange circles), with explored areas visually highlighted. Free cells adjacent to the target are marked pink and serve as detectable clues for the robots. Robots move at different speeds: explorer robots traverse a cell in one time step, while the rescuer robot takes two time steps for actions like moving forward or turning.
Figure 4: A snapshot of a simulated episode in the Minigrid environment, where two explorer agents (yellow and blue triangles) and one rescuer agent (red triangle) navigate an unknown area in search of a target (green square). The pink squares are detectable as clues to the target’s location. Agents’ current navigation goals are marked by corresponding colored circles.
6.2 Reward Structure
The main objective is for a rescuer robot to reach the target as quickly as possible. Each agent $i\in I$ receives a time-dependent reward $r^{i}_{success}(t)=300(1-0.9\frac{t}{h})$ upon the mission’s completion, incentivizing strategies that result in a quick rescue. Additionally, a similar time-dependent team reward $r^{i}_{locate}(t)=100(1-0.9\frac{t}{h})$ is given to all agents at the time step the target is discovered. To conserve energy and reduce unnecessary movement, agents incur a small penalty of $r^{i}_{movement}(t)=-0.05$ whenever they move to a new cell.
6.3 Macro-Action and Macro-Observation Spaces
Global goal candidates, or macro-actions, are defined as cells within a square of side length $L$ centered on the agent, giving rise to a discrete macro-action space of size $L^{2}$. Once a macro-action is selected, it is mapped to its corresponding coordinate $(x,y)$ on the grid map, establishing the agent’s global goal. Invalid actions, such as those targeting occupied or out-of-bounds cells, are handled using invalid action masking [35, 36].
Each agent’s macro-observation is composed of three elements: a global information map of size $S\times S\times 7$, a local information map of size $L\times L\times 6$, and a one-hot encoded agent class identifier. The channels in the local map provide information about cells in the agent’s immediate vicinity. These channels indicate which cells have been explored, the cells’ occupancy, the location of the target and clues around it, the agent’s current navigation goal, the locations of other rescuer robots, and the location of other explorer robots. Similarly, the global map represents this information for the entire map, with an extra channel representing the agent’s current location.
During centralized training or perfect communication conditions, these channels contain the latest information from other agents. However, during distributed execution when there is no communication link between explorer agent $i$ and rescuer agent $j$ for example, current location of agent $j$ would not show up in agent $i$’s rescuer agents’ location channel and agent $i$’s location would not be visible in agent $j$’s explorer agents’ location channel.
In Minigrid, the primitive action space includes four actions: moving forward, turning right, turning left, and stopping. An agent’s primitive actions are selected to follow the shortest path to the agent’s current navigation goal generated by the A* algorithm.
7 Simulation Results
We evaluated the effectiveness of the proposed CATMiP framework in the Minigrid simulation environment under varying environmental conditions, team compositions, and communication constraints. To assess CATMiP’s performance, we compared it against both multi-agent and single-agent planning-based exploration methods adapted for multi-agent scenarios.
7.1 Training Setup
Asynchronous Training: CATMiP was trained using asynchronous centralized training on an NVIDIA GeForce RTX™ 3090 GPU. The training was conducted across 64 parallel environments for 62,500 episodes with an episode horizon of 200 steps. Each episode included 3 agents: one rescuer agent, one explorer agent, and a third agent randomly assigned to either class. Macro-actions had a maximum duration of 10 time steps, after which agents automatically transitioned to standby mode. The asynchronous training process took approximately 285 hours.
Synchronous Training (Synch-CATMiP): For comparison, we trained a synchronous version of CATMiP, where macro-actions were updated for all agents simultaneously every 10 time steps. This synchronous training approach follows a procedure similar to MAT [26]. Due to fewer trajectories being collected and stored in the training buffer, training over the same total number of episodes required 166 hours.
In both cases, key hyperparameters were set as follows: embedding size $d=192$, local decision-making range $L=7$, global pooling range $G=4$, discount factor $\gamma=1$, a linearly decaying learning rate starting from $10^{-4}$, and using 10 training epochs. Both models have 1,280,786 total trainable parameters. Fig. 5 shows the progression of the average mission success rate and the average agent reward throughout the training for both models, where an exponential moving average with a span of 200 was applied for visualization. It can be seen that both models follow a similar trend in training for 62,500 episodes.
Figure 5: (a) Progression of the mission success rate and (b) progression of average agent rewards against the number of episodes during training of the two models.
7.2 Baseline Methods
We compare our method with planning-based exploration methods, including the multi-agent method of artificial potential field (APF) [37] and three single-agent frontier-based methods, namely a utility-maximizing algorithm (Utility) [38], a search-based nearest frontier method (Nearest) [39], and a rapid-exploring-random-tree-based method (RRT) [40]. The single-agent methods are adapted to multi-agent settings by planning on the shared global map, and their implementation on the Minigrid environment is adapted from Yu et al. [20]. During communication dropouts, the shared global map contains only the latest information received from other agents. All agent classes were treated identically during exploration. However, once a rescuer agent detects the target’s location, it immediately navigates along the shortest path to the target.
7.3 Evaluation Results and Analysis
The trained CATMiP models (both asynchronous and synchronous) and the planning-based baselines were evaluated on three tasks with increasing complexity. Task 1 involved one rescuer and one explorer agent in a $15\times 15$ grid. Task 2 increased the map size to $20\times 20$, with an additional explorer agent in the team. Task 3 scaled up to a $32\times 32$ environment with 6 agents consisting of 2 rescuers and 4 explorers. Evaluations are performed over 100 different episodes with the same random seed. Agents acted asynchronously during all evaluations.
Figure 6 shows the success rate of the different methods against mission time in Task 1. This comparison is made for three cases with different communication constraints. In Figure 6 (a) the agents have consistent communication throughout the mission, whereas in (b) and (c) the value of $\sigma$ in Equation 2 is set to $4$ and $2$ respectively, indicating increasing levels of communication loss. The learning-based models outperform the baselines in all three cases by showing higher success rates within the same time-frame. For example, as seen in Figure 6 (a), $97\%$ of the experiments using the CATMiP and Synch-CATMiP model were successful by the 100th step since the start of the mission, while the best performing baseline method, RRT, achieves $86\%$ success by the same time. Since the models were trained for a more complex task, this shows the scalibility of CATMiP to smaller environments and team sizes.
The same comparison between the models with different communication constraints is shown for Task 2 in Figure 7. Once again, the learning-based models outperform the baselines in all three communication scenarios.
For Task 3, with a map size of $32\times 32$ and $6$ agents, the learning-based models still show top performance alongside the planning-based methods RRT and Nearest, as shown in Figure 8. As the communication loss increases, CATMiP’s superiority becomes more prominent. Since CATMiP and Synch-CATMiP were trained on a smaller map of size $20\times 20$ and with a smaller team of agents, results on Task 3 show the scalability of our proposed framework to more complex tasks, with larger maps and team sizes. Moreover, in both Tasks 2 and 3, the CATMiP model which was trained with the asynchronous training scheme shows better adaptability to asynchronous execution than Synch-CATMiP, especially as the communication loss increases and agents have to act mostly based on their own observations only.
Figure 6: Success rate of different models on Task 1, with different communication constraints: (a) distributed execution with no communication loss, (b) distributed execution with moderate communication loss, and (c) distributed execution with heavy communication loss.
Figure 7: Success rate of different models on Task 2, with different communication constraints: (a) distributed execution with no communication loss, (b) distributed execution with moderate communication loss, and (c) distributed execution with heavy communication loss.
Figure 8: Success rate of different models on Task 3, with different communication constraints: (a) distributed execution with no communication loss, (b) distributed execution with moderate communication loss, and (c) distributed execution with heavy communication loss.
8 Conclusions & Future Work
This paper introduced CATMiP, a novel framework for coordinating heterogeneous multi-robot teams in environments with communication constraints. The proposed CMacDec-POMDP model provides the mathematical foundations of asynchronous and decentralized decision-making of heterogeneous agents by incorporating class-based distinctions across its components. Leveraging the transformer’s ability to handle variable input sequence lengths, the Multi-Agent Transformer architecture was extended to develop scalable coordination strategies that can be executed in a distributed manner by any number of agents.
CATMiP demonstrated robustness and scalability across different team sizes, compositions, and environmental complexities in simulations of search and target acquisition tasks. The results highlighted the framework’s adaptability to sporadic communication, asynchronous operations, and varied map conditions, achieving high mission success rates and competitive performance even under strict communication constraints. By addressing key challenges such as communication dropout, asynchronous operations, and agent heterogeneity, CATMiP shows significant potential for real-world applications where different types of mobile robots must cooperate under resource limitations and unpredictable conditions.
Future research will investigate integrating temporal memory into the network and expanding the framework to scenarios involving dynamic targets. The system’s performance will also be evaluated in more realistic 3D simulation environments as well as real-world experiments.